# Unconventional criticality in $O(D)$-invariant loop-constrained Landau theory

### v2

@@ Abstract
We study an unconventional phase transition in ferroelectrics where the polarization field is constrained to be divergence-free, allowing only loop-like configurations. This local constraint fundamentally alters the critical behavior, driving the system beyond the Landau-Ginzburg-Wilson paradigm. A renormalization group analysis shows that the polarization acquires an unusually large anomalous dimension, $\eta\approx 0.239$ in three dimensions, far exceeding the typical values in $O(3)$-invariant systems. We attribute this effect to an emergent gauge symmetry originating from the zero divergence constraint. Such gauge-field behavior is reminiscent of fractionalized phases, revealing a fundamental connection between constrained ferroelectrics and emergent gauge phenomena in correlated matter.
@@
These are just a few examples of the diverse classical and quantum systems where divergence-free conditions govern the low-energy physics, generating gauge structures and long-range correlations. When supplemented with continuous symmetries such as the orthogonal group $O(D)$, where $D$ is the spatial dimension, these systems can exhibit LGW-incompatible critical behavior, which is in the focus of this Letter. As an exemplary system revealing the constraint-imposed effects, we consider ferroelectric materials where the polarization is divergence-free, ${\mbox{$\nabla$}}\cdot\bm{P}=0$, giving rise to topological configurations that influence the system’s critical properties. We refer to such materials as topological ferroelectrics.
@@
A natural question emerging from the above considerations concerns the role played by the zero divergence constraint in determining the phase structure of the physical system. How much does such a constraint affect the LGW theory of a ferroelectric system? Remarkably, we find that the critical behavior of correlation functions is affected significantly, reflecting a scenario that is usually encountered in physical systems exhibiting fractionalization. The latter is a property underlying the decay of an order parameter into emergent building blocks, similarly to the physics of quark deconfinement as studied in high-energy physics. In condensed matter physics, it is typically associated to an emerging gauge symmetry, with realizations in quantum antiferromagnets featuring competing orders. This typically leads to a phase structure that is incompatible with the LGW paradigm [1, 4]. The corresponding critical behavior where competing order parameters decay into fractionalized quasi-particles is what one generally refers to DQC. One of the most salient features underlying this scenario is that it leads to a unusually large anomalous dimension, $\eta$, of the order parameter $O(\bm{x})$ [1]. Recall that in LGW-compatible systems $\eta$ is typically considerably smaller than the thermodynamic critical exponents, namely, those associated to the critical singularities of the free energy density.
@@
In this Letter, we show that accounting for the constraint ${\mbox{$\nabla$}}\cdot\bm{P}=0$ implies a large anomalous dimension, with no obvious presence of fractionalized excitations. In fact, we find at two-loop order in perturbation theory that $\eta=0.239$ in three dimensions. By contrast, in a LGW theory with $O(N)$ symmetry one finds a value of $\eta$ which is typically one order of magnitude smaller than the value found here. Indeed, in an $O(3)$ symmetric theory, which is the natural case to compare with in the present context, the result is $\eta\approx 0.034$. Crucially, the constraint ${\mbox{$\nabla$}}\cdot\bm{P}=0$ naturally imposes that $N=D$, i.e., the number of components of the order parameter coincides with spatial dimensionality.
@@
Constraint-generated loops – The zero divergence constraint in topological ferroelectrics can be viewed from a statistical field theory perspective as representing the collective behavior of objects that we call here polarization loops, or $P$-loops, which have many aspects in common with vortex loops in three-dimensional superfluids. At this point it is worth mentioning the well known dielectric approach to understand phase transitions in two-dimensional superfluids [25]. While three-dimensional superfluids feature vortex loop excitations [18, 19, 20], in two dimensions vortices are point-like objects that arise in pairs having opposite vorticities (++ or -[24]. This is more easily seen by observing that a vortex loop pierces a plane at two points. The point where the loop flows to corresponds to (say) positive vorticity, the one where it outflows yields a negative vorticity. Hence, an ensemble of point vortices in a two-dimensional superfluid can be viewed as a Coulomb gas [26] where the vortices represent electric charges. Neutrality of the Coulomb gas is guaranteed by the fact that point vortices arise only in pairs of opposite vorticities [25]. A two-dimensional Coulomb gas undergoes a phase transition from a dielectric phase where vortex pairs are bound into dipoles to a plasma phase where the dipoles unbind. This is the physical picture behind the BKT phase transition. In the context of this Letter, such a Coulomb gas is actually a dual representation of the two-dimensional $O(2)$ nonlinear $\sigma$-model ($N=D=2$). As a model for a two-dimensional superfluid, the latter is also subject to a zero divergence constraint, namely, the superfluid velocity $\bm{v}_{s}$ satisfies the incompressibility condition ${\mbox{$\nabla$}}\cdot\bm{v}_{s}=0$.
@@
does not vanish in general. In terms of vortex loops, the polarization density is just the circulation of the vorticity, in a way reminiscent of the old string theory [33]. We should note that the vanishing of the total dipole moment does not occur when thermal fluctuations are accounted for, as we demonstrate further in this Letter.
@@
To study the critical properties of the presented $P$-loops, we employ an LGW type of theory where the constraint ${\mbox{$\nabla$}}\cdot\bm{P}=0$ is enforced, in a situation characteristic of topological ferroelectrics.
Renormalization group analysis – In the absence of anisotropies, the action of a topological ferroelectric is given in $D$ dimensions by,
@@
where $\bm{P}$ is a $D$-dimensional polarization vector, see Eq. (1), and the constraint ${\mbox{$\nabla$}}\cdot\bm{P}=0$ holds, arising from the absence of depolarization charges in the low-energy states. This implies that only closed-loop configurations of polarization corresponding to bound dipole textures are energetically favored and contribute to the partition function.
@@
For small $\hat{a}$ the theory features an infrared stable fixed point, but the universality class is clearly not the $O(D)$ one. Indeed, linearizing the RG equations (9) near the fixed point leads to the following thermal eigenvalue governing the behavior of the correlation length,
@@
Transverse correlation function – The zero divergence constraint of the polarization implies that the correlation function $\Delta_{ij}(\bm{r})=\langle P_{i}(\bm{r})P_{j}(0)\rangle$ is transverse, $\partial_{i}\Delta_{ij}(\bm{r})=0$, or, equivalently, $q_{i}\tilde{\Delta}_{ij}(\bm{q})=0$ in momentum space. By inverting the kernel $\left(q^{2}+a\right)\delta_{ij}-{q_{i}q_{j}}/{\alpha}$ and setting the parameter $\alpha\to 0$, the free propagator can be written as
@@
The calculation of the tadpole correction to this propagator (Appendix LABEL:app:Tadpole) shows that any fluctuation correction to the mass term is transverse. It also recovers the same RG equation for $\hat{a}$ in (9) as obtained using the effective action approach. At the next step, we calculate the two-loop correction required to estimate the anomalous dimension $\eta$ of a system under divergence-free conditions.
@@
$\mathcal{L}_{\rm int}=\frac{1}{8}b_{kmnp}P_{k}P_{m}P_{n}P_{p}.$ (12)

### v3

@@ Abstract
We study an unconventional phase transition in ferroelectrics where the polarization field is constrained to be divergence-free, allowing only loop-like configurations. This local constraint fundamentally alters the critical behavior, driving the system beyond the Landau-Ginzburg-Wilson paradigm. A renormalization group analysis shows that the polarization acquires an unusually large anomalous dimension, $\eta\approx 0.239$ in three dimensions, far exceeding the typical values in $O(3)$-invariant systems. We attribute this effect to a naturally induced gauge symmetry originating from the zero divergence constraint. Such constraint-induced gauge-field behavior is reminiscent of fractionalized phases, revealing a fundamental connection between constrained ferroelectrics and emergent gauge phenomena in correlated matter.
@@
These are just a few examples of the diverse classical and quantum systems where divergence-free conditions govern the low-energy physics, generating gauge structures and long-range correlations. When supplemented with continuous symmetries such as the orthogonal group $O(D)$, where $D$ is the spatial dimension, these systems can exhibit LGW-incompatible critical behavior, which is the focus of this work. Crucially, we address the role of the functional measure of the partition function as a significant factor altering the universality class of the system. This means that the phase structure of the system is not only determined by the action and its symmetries. As an exemplary system revealing such a behavior, we analyze ferroelectric materials where the polarization is constrained to be divergence-free, ${\mbox{$\nabla$}}\cdot\bm{P}=0$, giving rise to topological configurations that influence the system’s critical properties. We refer to such materials as topological ferroelectrics. This constraint is precisely accounted for in the integration measure of the partition function.
@@
A natural question emerging from the above considerations concerns the role played by the zero divergence constraint in determining the phase structure of the physical system. How much does such a constraint affect the LGW theory of a ferroelectric system? Remarkably, we find that the critical behavior of correlation functions is affected significantly, reflecting a scenario that is usually encountered in physical systems exhibiting fractionalization. The latter is a property underlying the decay of an order parameter into emergent building blocks, similarly to the physics of quark deconfinement as studied in high-energy physics. In condensed matter physics, this is generally associated to an emergent gauge symmetry, with realizations in quantum antiferromagnets featuring competing orders. This leads to a phase structure incompatible with the LGW paradigm [1, 4]. One of the most salient features underlying this scenario is that it leads to a unusually large anomalous dimension, $\eta$, of the order parameter $O(\bm{x})$ [1]. Recall that LGW-compatible systems have considerably smaller values of $\eta$ compared to those of thermodynamic critical exponents, namely, exponents obtained from the scaling of the free energy density [25, 26].
@@
The route towards non-LGW behavior in this work differs substantially from the DQC mechanism, as it does not involve fractionalization and keeps the LGW action intact. However, in spite of this, the partition function is changed in an essential way due to a constrained integral measure on the fluctuations. This in turn considerably affects the critical behavior of the order parameter correlation functions, leading to an unusually large anomalous dimension, similar to DQC. Specifically, we show that accounting for the constraint ${\mbox{$\nabla$}}\cdot\bm{P}=0$ for the electric polarization indeed implies a large anomalous dimension without apparent fractionalized excitations. In fact, we find at two-loop order in perturbation theory that $\eta=0.239$ in three dimensions. By contrast, in a LGW theory with $O(N)$ symmetry one finds a value of $\eta$ which is typically one order of magnitude smaller than the value found here. Indeed, in an $O(3)$ symmetric theory, which is the natural case to compare with in the present context, the result is $\eta\approx 0.034$. Crucially, the constraint ${\mbox{$\nabla$}}\cdot\bm{P}=0$ naturally imposes that $N=D$, i.e., the number of components of the order parameter coincides with spatial dimensionality.
@@
Constraint-generated loops – The zero divergence constraint in topological ferroelectrics can be viewed from a statistical field theory perspective as representing the collective behavior of objects that we call here polarization loops, or $P$-loops, which have many aspects in common with vortex loops in three-dimensional superfluids. At this point it is worth mentioning the well known dielectric approach to understand phase transitions in two-dimensional superfluids [27]. While three-dimensional superfluids feature vortex loop excitations [18, 19, 20], in two dimensions vortices are point-like objects that arise in pairs having opposite vorticities (++ or -[24]. This is more easily seen by observing that a vortex loop pierces a plane at two points. The point where the loop flows to corresponds to (say) positive vorticity, the one where it outflows yields a negative vorticity. Hence, an ensemble of point vortices in a two-dimensional superfluid can be viewed as a Coulomb gas [28] where the vortices represent electric charges. Neutrality of the Coulomb gas is guaranteed by the fact that point vortices arise only in pairs of opposite vorticities [27]. A two-dimensional Coulomb gas undergoes a phase transition from a dielectric phase where vortex pairs are bound into dipoles to a plasma phase where the dipoles unbind. This is the physical picture behind the BKT phase transition. In the context of this work, such a Coulomb gas is actually a dual representation of the two-dimensional $O(2)$ nonlinear $\sigma$-model ($N=D=2$). As a model for a two-dimensional superfluid, the latter is also subject to a zero divergence constraint, namely, the superfluid velocity $\bm{v}_{s}$ satisfies the incompressibility condition ${\mbox{$\nabla$}}\cdot\bm{v}_{s}=0$.
@@
does not vanish in general. We should note that the vanishing of the total dipole moment does not occur when thermal fluctuations are accounted for, as we demonstrate further in this work.
@@
To study the critical properties of the presented $P$-loops, we employ an LGW type of theory where the constraint ${\mbox{$\nabla$}}\cdot\bm{P}=0$ is enforced, in a situation characteristic of topological ferroelectrics. Importantly, although our model is LGW-based, the divergence-free constraint removes the standard assumption that the order parameter components fluctuate independently and subjected only to the internal symmetries of the action. In the case considered here the constraint must be incorporated in the integral measure of the partition function. This places the description beyond the LGW paradigm, because in the study of critical fluctuations it is the partition function rather than the action that matters. The latter still has a LGW form though, but integrating over the fields in the partition function requires changing the integral measure to accommodate the local constraint ${\mbox{$\nabla$}}\cdot\bm{P}=0$.
Renormalization group analysis – In the absence of anisotropies, the action of a topological ferroelectric is given in $D$ dimensions by [35],
@@
where $\bm{P}$ is a $D$-dimensional polarization vector, see Eq. (1), and the constraint ${\mbox{$\nabla$}}\cdot\bm{P}=0$ holds, arising from the absence of depolarization charges in the low-energy states. This implies that only closed-loop configurations of polarization corresponding to bound dipole textures are energetically favored and contribute to the partition function. Recall that in dielectrics the polarization couples linearly to the electric field, which implies that the “mass” coefficient $a$ appearing in the action corresponds to the inverse susceptibility. Ferroelectric order, on the other hand, causes the electric field to depend nonlinearly in the polarization, so higher powers of $\bm{P}$ arise in the expression of the electric field, and ferroelectricity is induced by the sign change in $a$ as the temperature is reduced below the transition temperature, thus driving the onset of a spontaneous polarization. The coefficient $b>0$ physically controls the stability of the energy in the spontaneous symmetry breaking process [35].
@@
For small $\hat{a}$ the theory features an infrared stable fixed point, but the universality class is clearly not the $O(D)$ one. Indeed, linearizing the RG equations (9) near the fixed point leads to the following thermal eigenvalue, i.e. the RG eigenvalue associated with the relevant direction that tunes the system to criticality, governing the behavior of the correlation length,
@@
Transverse correlation function – The zero divergence constraint of the polarization implies that the correlation function $\Delta_{ij}(\bm{r})=\langle P_{i}(\bm{r})P_{j}(0)\rangle$ is transverse, $\partial_{i}\Delta_{ij}(\bm{r})=0$, or, equivalently, $q_{i}\tilde{\Delta}_{ij}(\bm{q})=0$ in momentum space. To calculate it, we introduce a quadratic term $\alpha\lambda^{2}/2$ for the auxiliary field $\lambda$ in Eq. (6). Integrating out $\lambda$ generates the term $({\mbox{$\nabla$}}\cdot\bm{P})^{2}/(2\alpha)$ in the effective action. Hence, by inverting the kernel $\left(q^{2}+a\right)\delta_{ij}+{q_{i}q_{j}}/{\alpha}$ and setting the tuning parameter $\alpha\to 0$ (see Appendix LABEL:app:no_constraint), the free propagator becomes,
@@
The corresponding free propagator in real space, $\Delta_{ij}^{\left(0\right)}\left(\bm{r}\right)$, is given in Appendix LABEL:app:PropReal. The calculation of the tadpole correction to the propagator (11), see Appendix LABEL:app:Tadpole, shows that any fluctuation correction to the mass term is transverse. It also recovers the same RG equation for $\hat{a}$ in (9) as obtained using the effective action approach.
At the next step, we calculate the two-loop correction required to estimate the anomalous dimension $\eta$ of a system under divergence-free conditions.
@@
$\mathcal{L}_{\rm int}=\frac{1}{8}b_{kmnp}P_{k}P_{m}P_{n}P_{p},$ (12)
where, for an isotropic system,
Performing the calculations for the two-point function to order $b^{2}$ (see Appendix LABEL:app:Sunrise) yields for the two-loop propagator in $D=3$,
where $C_{1}=11/6$, $C_{2}\approx 0.04$, and $c$ contains the UV cutoff of the real-space integrals, i.e. $c\to 0$.
Next, we rewrite the propagator (14) in the form
where $Z$ is the field renormalization constant, and $\eta$ is calculated at the fixed point. More precisely, the process of renormalization involves the replacement of bare quantities by renormalized ones, which are overall independent of $\Lambda$ and are given as a perturbative expansion in terms of the bare parameters of the theory. The requirement that all renormalized correlation functions are $\Lambda$-independent yields the RG functions [26]. In the case of $\eta$, we have explicitly (see Appendix LABEL:app:EtaDerive),
Conclusions and final remarks – We have shown that imposing a local divergence-free constraint on the polarization field drives a ferroelectric system into a universality class that lies beyond the conventional LGW framework. This profoundly modifies the configuration space, restricting it to loop-like polarization structures and thereby altering the critical behavior. The RG analysis demonstrates that the polarization correlation function features a large anomalous dimension in three dimensions, $\eta=0.239$, which significantly exceeds that of ordinary $O(3)$ LGW models.
This unconventional scaling originates from a naturally induced $U(1)$ gauge symmetry, which becomes explicit when the constraint is resolved as $\bm{P}={\mbox{$\nabla$}}\times\bm{A}$, rendering the effective theory manifestly gauge invariant. In this sense, the zero divergence constraint is understood as a local conservation law corresponding to a local gauge symmetry. This leads to the emergence of a gauge field and therefore to a large value of $\eta$. Usually, the argument runs the opposite way, with a gauge symmetry yielding a local conservation law, one example being current conservation in electrodynamics. From a field theoretic perspective, at the critical point current conservation implies that currents do not have an anomalous dimension [41]. This fact is a consequence of the scale invariance of the critical point. The opposite is true here: the order parameter, viewed as a conserved quantity, has a large anomalous dimension.
Acknowledgements
Data availability
No data were created or analyzed in this study.
Appendices
Appendix A: Arnold’s helicity theorem – The helicity, or Hopf invariant, of a divergence-free field $\bm{P}$ defined in a volume $V$ can be expressed through its divergenceless vector potential $\bm{A}$, i.e. $\bm{P}={\mbox{$\nabla$}}\times\bm{A}$, ${\mbox{$\nabla$}}\cdot\bm{A}=0$, as [10]
$\mathscr{H}=\int_{V}d^{3}r\,\bm{P}\cdot\bm{A}.$ (18)
The $P$-loop representation of polarization vector (1) assumes the vector potential $\bm{A}(\bm{r})$,
$\bm{A}$ $=$ $-\int_{V}d^{3}r^{\prime}\frac{{\mbox{$\nabla$}}^{\prime}\times\bm{P}(\bm{r^{\prime}})}{4\pi\lvert\bm{r}-\bm{r^{\prime}}\rvert}=\int_{V}d^{3}r^{\prime}\frac{\left(\bm{r}-\bm{r^{\prime}}\right)\times\bm{P}(\bm{r^{\prime}})}{4\pi\lvert\bm{r}-\bm{r^{\prime}}\rvert^{3}}$ (19)
$=$ $\frac{1}{4\pi}\sum_{a}q_{a}\int_{V}d^{3}r^{\prime}\frac{\left(\bm{r}-\bm{r^{\prime}}\right)\times\oint_{L_{a}}d{\mbox{$\gamma$}}^{(a)}\delta^{3}\left(\bm{r^{\prime}}-{\mbox{$\gamma$}}^{(a)}\right)}{\lvert\bm{r}-\bm{r^{\prime}}\rvert^{3}}$
$=$ $-\frac{1}{4\pi}\sum_{a}q_{a}\oint_{L_{a}}\frac{d{\mbox{$\gamma$}}^{(a)}\times\left(\bm{r}-{\mbox{$\gamma$}}^{(a)}\right)}{\lvert\bm{r}-{\mbox{$\gamma$}}^{(a)}\rvert^{3}}.$
Then, the helicity (18) becomes
$\mathscr{H}$ $=$ $-\frac{1}{4\pi}\int_{V}d^{3}r\,\sum_{b}q_{b}\oint_{L_{b}}d{\mbox{$\gamma$}}^{(b)}\delta^{3}\left(\bm{r}-{\mbox{$\gamma$}}^{(b)}\right)$
$\cdot\sum_{a}q_{a}\oint_{L_{a}}\frac{d{\mbox{$\gamma$}}^{(a)}\times\left(\bm{r}-{\mbox{$\gamma$}}^{(a)}\right)}{\lvert\bm{r}-{\mbox{$\gamma$}}^{(a)}\rvert^{3}}$
$=$ $\frac{1}{4\pi}\sum_{a,b}q_{a}q_{b}\oint\limits_{L_{a}}\oint\limits_{L_{b}}\frac{\left[d{\mbox{$\gamma$}}^{(b)}\times d{\mbox{$\gamma$}}^{(a)}\right]\cdot\left({\mbox{$\gamma$}}^{(b)}-{\mbox{$\gamma$}}^{(a)}\right)}{\lvert{\mbox{$\gamma$}}^{(b)}-{\mbox{$\gamma$}}^{(a)}\rvert^{3}}.$
Hence, the helicity is expressed in terms of the Gauss linking integral. It represents the linking number of the curves $L_{a}$ and $L_{b}$, and corresponds to the average linking of the polarization field streamlines [14], $\mathscr{L}_{\bm{P}}=\mathscr{H}$.
Appendix B: Renormalization group calculations – Integrating out the Gaussian fluctuations of the polarization $\widetilde{P}_{j}$ in (7) yields the following effective action,
$S_{\rm eff}$ $=$ $V\left(\frac{a}{2}P_{0}^{2}+\frac{b}{8}P_{0}^{4}\right)+\frac{1}{2}{\rm Tr}\ln\left(-\nabla^{2}+a+\frac{3}{2}bP_{0}^{2}\right)$ (21)
$+$ $\frac{D-1}{2}{\rm Tr}\ln\left(-\nabla^{2}+a+\frac{bP_{0}^{2}}{2}\right)$
[…]
$+$ $\left.\sum_{j=2}^{D}\partial_{j}\lambda(\bm{r})G_{T}(\bm{r}-\bm{r}^{\prime})\partial_{j}^{\prime}\lambda(\bm{r}^{\prime})\right],$
where $V$ is the (infinite) volume, and $G_{L}(\bm{r}-\bm{r}^{\prime})$ and $G_{T}(\bm{r}-\bm{r}^{\prime})$ have Fourier transforms,
$G_{L}(\bm{q})=\frac{1}{q^{2}+a+\frac{3bP_{0}^{2}}{2}},\quad G_{T}(\bm{q})=\frac{1}{q^{2}+a+\frac{bP_{0}^{2}}{2}}.$ (22)
Next, we integrate out the Lagrangian multiplier field, which gives the following effective action $s_{\rm eff}$ per unit volume (which in this case corresponds to the free energy density) written in terms of momentum space variables,
$s_{\rm eff}$ $=$ $\frac{a}{2}P_{0}^{2}+\frac{b}{8}P_{0}^{4}+\left(\frac{D}{2}-1\right)\int_{q}\ln\left(q^{2}+a+\frac{bP_{0}^{2}}{2}\right)$ (23)
$+$ $\frac{1}{2}\int_{q}\ln\left[q^{2}\left(q^{2}+a+\frac{bP_{0}^{2}}{2}\right)+bP_{0}^{2}\sum_{j=2}^{D}q_{j}^{2}\right],$
where $\int_{q}\equiv\int\frac{d^{D}q}{(2\pi)^{D}}$. The above form of the effective action reveals a different pattern of would-be Goldstone modes. Typically, without the constraint ${\mbox{$\nabla$}}\cdot\bm{P}=0$, the minimum of the effective potential in the spontaneous symmetry breaking regime would yield D1D-1 gapless transverse modes and one longitudinal gapped mode, as required by the $O(D)$ symmetry. At one loop, the minimum is given by the tree level contribution, $P_{0}^{2}=-2a/b$, requiring $a<0$. From Eq. (23), all modes are gapless as $q\to 0$, a direct consequence of the constraint, which imposes transversality of $\bm{P}$. A gapped mode and the usual Goldstone counting appear only if $q_{1}\to 0$ faster than the other $\bm{q}$ components, reminiscent of early studies of the critical behavior in systems with dipolar interaction [42].
In order to obtain the one-loop RG equations, we expand $s_{\rm eff}$ up to the second order in $b$, which results in the expression (Unconventional criticality in $O(D)$-invariant loop-constrained Landau theory) for the effective action.
Appendix C: RG functions with no constraint – In the absence of the ${\mbox{$\nabla$}}\cdot\bm{P}=0$ constraint, the one-loop RG equations (9) would acquire the usual $O(N)$ form,
$\Lambda\frac{d\hat{a}}{d\Lambda}=-2\hat{a}-\frac{N+2}{2}\frac{\hat{b}}{1+\hat{a}}$, (24)
$\Lambda\frac{d\hat{b}}{d\Lambda}=-(4-D)\hat{b}+\frac{N+8}{2}\frac{\hat{b}^{2}}{(1+\hat{a})^{2}}.$
It is clearly seen that imposing $N=D$ directly does not lead to Eqs. (9), which highlights the nontrivial impact of the constraint on the critical properties of the system.
To illustrate the effect of the constraint, one may keep the tuning parameter $\alpha$ in the free propagator (11) as:
$\tilde{\Delta}_{ij}^{(0)}(\bm{q}\,;\alpha)=\frac{1}{q^{2}+a}\left(\delta_{ij}-\frac{q_{i}q_{j}}{(1+\alpha)q^{2}+\alpha a}\right),$ (25)
thus setting $\alpha\to 0$ results in the considered constrained case, while $\alpha\to\infty$ corresponds to the absence of the constraint. The $\alpha$-dependent RG $\beta$-functions read
$\Lambda\frac{d\hat{a}}{d\Lambda}$ $=$ $-2\hat{a}-\frac{1}{2}\left(D-\frac{2}{D}+1\right)\frac{\hat{b}}{1+\hat{a}}-\frac{D+2}{2D}\frac{\frac{\alpha}{1+\alpha}\hat{b}}{1+\frac{\alpha}{1+\alpha}\hat{a}}$,
$\Lambda\frac{d\hat{b}}{d\Lambda}$ $=$ $-(4-D)\hat{b}+\left[\frac{D+7}{2}-\frac{6(D+1)}{D(D+2)}\right]\frac{\hat{b}^{2}}{(1+\hat{a})^{2}}$
$+$ $\frac{\alpha}{1+\alpha}\frac{4(D-1)}{D(D+2)}\frac{\hat{b}^{2}}{(1+\hat{a})(1+\frac{\alpha}{1+\alpha}\hat{a})}$
$+$ $\frac{1}{2}\frac{\alpha^{2}}{(1+\alpha)^{2}}\left[1+\frac{4}{D}+\frac{12}{D(D+2)}\right]\frac{\hat{b}^{2}}{(1+\frac{\alpha}{1+\alpha}\hat{a})^{2}},$
and reduce to Eqs. (9) and (24) with $N=D$ in two limit cases $\alpha\to 0$ and $\alpha\to\infty$, respectively.
Appendix D: Free polarization proparator in real space – The constraint ${\mbox{$\nabla$}}\cdot\bm{P}=0$ imposes that in the absence of interactions the polarization propagator is given in momentum space by Eq. (11). To obtain the corresponding propagator in real space, we rewrite the propagator as,
and introduce the function,
$C(r;a)$ $=$ $\int\frac{d^{D}q}{(2\pi)^{D}}\frac{e^{i\bm{q}\cdot\bm{r}}}{q^{2}+a}$ (28)
$=$ $\frac{1}{(2\pi)^{D/2}}\left(\frac{\sqrt{a}}{r}\right)^{\frac{D-2}{2}}K_{\frac{D-2}{2}}(\sqrt{a}r),$
which for $D=3$ holds exactly. Then, in real space,
$=$ $\frac{1}{4\pi ar^{D}}\left(\frac{1}{2\pi}\right)^{\frac{D-3}{2}}$
[…]
For $D=3$, this expression simplifies to
allowing us to calculate the two-loop propagator (14).
Appendix E: Tadpole correction to the propagator – The one-loop propagator is given by,
$=$ $\frac{1}{q^{2}+a}\left(\delta_{in}-\frac{q_{i}q_{n}}{q^{2}}\right)$
[…]
where $\int_{k}\equiv\int\frac{d^{D}k}{(2\pi)^{D}}$ and